### v1

@@ Abstract
Conformal phased arrays provide multiple degrees of freedom to the scan angle, which is typically limited by antenna aperture in rigid arrays. Silicon-based RF signal processing offers reliable, reconfigurable, multi-functional, and compact control for conformal phased arrays that can be used for on-the-move communication. While the lightweight, compactness, and shape-changing properties of the conformal phased arrays are attractive, these features result in dynamic deformation of the array during motion leading to significant dynamic beam pointing errors. We propose a silicon-based, compact, reconfigurable solution to self-correct these dynamic deformation-induced beam pointing errors. Furthermore, additive printing is leveraged to enhance the flexibility of the conformal phased arrays, as the printed conductive ink is more flexible than bulk copper and can be easily deposited on flexible sheets using different printing tools, providing an environmentally-friendly solution for large-scale production. The inks such as conventional silver inks are expensive and copper-based printable inks suffer from spontaneous metal oxidation that alters trace impedance and degrades beamforming performance. This work uses a low-cost molecular copper decomposition ink with reliable RF properties at different temperature and strain to print the proposed intelligent conformal phased array operating at 2.1 GHz. Proof-of-concept prototype $2\times 2$ array self-corrects the deformation induces beampointing error with an error $<1.25^{\circ}$. The silicon based array processing part occupying only 2.56 mm2 area and 78.5 mW power per tile.
@@
Additive Printing, conformal phased arrays, deformation correction, molecular copper decomposition ink, self-correction loop, scalable tile-based.
@@ 1 Introduction
Phased array systems support directional data transfer which enable inherent spatial selection to support ever increasing wireless data rate. Initial research on the phased arrays started several decades back (earliest reported in 1909) to enable inertia-free fast beam steering by avoiding mechanical gimballing [1, 2]. The underlying assumption in these works was that the antenna surfaces were rigid. By the 1940’s, conformal arrays were developed to curve around spherical and nonuniform surfaces [3, 4, 5, 6] that afforded array systems an added degree of freedom by eliminating the conventional radome. Conformal phased arrays have been used extensively since then in the automotive, aviation, and space industries. Having non-uniform antenna structure enhances the Field-of-View (FoV). When deployed on an aircraft, the array benefits from the aerodynamic shape of the plane which improves the fidelity of the in-flight wireless and makes the airborne systems such as Unmanned Aerial Vehicles (UAV) management easy as shown in Fig. 1.
@@
Satellite communication (SATCOM) has embraced conformal phased arrays due to their lightweight, small volume, and shape-changing properties. The manageable costs associated with launching and deploying[7, 8] have led to the widespread adoption of conformal phased arrays in Low-Earth Orbit (LEO) satellites[9]. LEO satellite networks such as Amazon Kuiper  [10], SpaceX Starlink [11], OneWeb, and Telesat Light-speed, boast coverage fields that collectively span the entire terrestrial surface area, opening profound opportunities for direct-to-consumer wireless communications in remote areas [12, 13]. These satellite networks, in conjunction with wearable/textile arrays, will further aid challenging research expeditions and facilitate disaster relief management as illustrated in Fig. 1.
Figure 1: Dynamic wing load and wing vibration deformation on in-flight wireless and Unmanned Aerial Vehicles (UAV). Proposed dynamic deformation correction technique using silicon based self-calibration loop.
@@
Surface deformation of a conformal phased array with curvature enhances its Field-of-View (FoV) but incurs increased inter-element coupling[14] and path length variations leading to beam pointing errors [15]. For satellites, it is possible to estimate the amount of deformation following the expansion of arrays once they are placed in orbit. While the initial deformation due to the aerodynamic shape of aircraft and drones can be estimated, the dynamic deformations arising from wing loading and vibrations[16] during transit cannot be predetermined. These dynamic deformations depend on factors such as array weight, size, drone/aircraft speed, wind conditions, initial deformation, etc. Intelligent self-adaptive conformal phased arrays will serve as the cornerstone to enable existing and emerging applications as shown in Fig. 1, while overcoming limitations from dynamic deformations [17].
@@
Recent studies have explored numerous strategies for correcting the radiation pattern of conformal phased arrays [18, 19, 20, 21] arising from physical deformities. However, element-level deformation sensing and compensation with conventional mechanical strain sensors [22, 23] and sensing mutual coupling [24, 25] has been ineffective due to complexity and non-planar shapes beyond single-point curvature. Machine-learning (ML), and in particular, pre-trained deep-learning based networks are increasingly being used for adaptive beam synthesis allowing arbitrary shape correction based on training [16]. However, it relies heavily on feature-rich longitudinal-series training data with prior knowledge of the deformation surface and uncertain operational environments. These models further require separate dictionary sets for near-field and far-field patterns as well as environmental conditions, demanding exponentially increasing storage as the array size is scaled. The multi-dimensional variable dynamic range search algorithm in [26], employs array-level compensation suited for both near- and far-field. However, its higher computational complexity and the additional requirement for separate generation and recovery units with high computing power along with low-power transceivers makes this technique challenging to adapt for airborne platoons and portable units. Phase correction using iterative methods such as genetic algorithm [27] and iterative phase synthesis [28, 29, 30] not only requires external engines for executing the iterative algorithm, but also incurs large latencies not suited for fast communications-on-the-move applications.
@@
Figure 2: 3D printed array layer stackup and self a. Full break down of Inks in each layers of array, Ball Milled Cu Ink b. SEM (Scanning Electron Microscope) of the ink before sintering, and after sintering c. Test setup for resistance measurements at different temperatures: Printed traces using the copper molecular decomposition ink are placed on a hot plate, and the temperature is varied to note the variation in resistance with respect to temperature. d. Measured resistance at different temperatures.
@@ 2.1 Low-cost reliable molecular copper decomposition ink for additive printing
Additively printed RF circuits have gained prominence in next-generation wireless systems with their ability to generate non-uniform conformal shapes required for large-scale infrastructure while being environmentally-friendly [31, 32, 33, 34, 35, 36, 37]. Recent works for additive printing techniques such as aerosol jetting, ink jetting, direct writing, and screen printing have shown enhanced scalability, reduced material waste generation, and cost-effective manufacturing processes. However, the efficacy of additive printing method is heavily contingent upon the quality and properties of the conductive inks utilized in the application. Common challenges in additive manufacturing of conductive inks revolve around printability and electrical performance [38]. Factors such as additives, ink viscosity, uniformity, micro/nano material structure, and size directly influence the potential printing method, substrate material, sintering conditions, and electrical performance of the traces. The electrical performance of printed conductive traces is also commonly evaluated for stability under repeated exposure to high temperatures, corrosive environments, and mechanical bending and torsion, reflecting long-term reliability during use.
@@
Figure 3: a. Dipole antenna printed using proposed Cu decomposition ink testing under varied strains b. Laser ablated dipole antenna tested at different temperatures on a hotplate c. Test set up for measuring ink stability under different strains. d. Measured S11 of the dipole antenna printed using Cu ink to demonstrate ink stability at different strains and temperatures.
Although silver (Ag)-based inks have been largely appreciated due to their low resistivity (1.59 $\mu\Omega.cm$) for use in printed electronics, their high-cost margins make copper (Cu) a much more cost-effective alternative with potential for comparable resistivity performance (1.72 $\mu\Omega.cm$). Excessive additives for printability and potential for ambient condition oxidation of Cu nanostructured inks also suggest a greater potential for industrial applications of molecular Cu-based inks. Herein a molecular Cu formate-based ink is utilized to create a highly conductive Cu thin-film ($=$35 MS/m) [40] on a Pyrallux substrate with demonstratively consistent electrical, and RF performance under varied strains and temperatures.
@@ 2.2 Scalable tile-based delay-phase tunable receiver array
Emergence of highly integrated silicon-based RF signal processors have revolutionized array synthesis and beam steering with multi-modal spatial signal processing [41, 42, 43, 44] enabling high signal-to-noise ratios in commercial arrays. When the number of elements increases, the beam width reduces, thereby enhancing the spatial selectivity. At 2.1 GHz, the antenna dimensions are in centimeters, and therefore scaling to a larger array is not straightforward. This scaling comes with several challenges [45], such as additive printing of high-resolution structures with physically large dimensions and combining signals from multiple antennas over wide bandwidth with high signal-to-noise ratio. We are adopting a modular tile-based approach for the receiver array as shown conceptually in the illustration in Fig. 2. In this work, each tile consists of a $2\times 2$ conformal antenna array attached to a 4-channel beamforming integrated circuit (BFIC). Figure 5e shows 4 tiles thermally bonded together to form a 16-element array. The BFIC in each tile combines each antenna signals and therefore act as a sub-array. Each sub-array outputs are then combined using a second-stage combiner to achieve $4\times 4$ radiation pattern. The gain, phase shiffter and impedance mismatches, and delay skews exacerbates as the array scales. The modular tile-based array structure thus enables easier calibration of larger arrays. An N-element array can be divided and processed into N/4 sub-arrays.
@@
The BFIC in a $2\times 2$ tile performs discrete-time true-time-delay BFIC as shown in Fig. 4a. The developed ink and the printing technique ensure that the four channels of the BFIC are connected to each antenna through impedance-matched via holes and traces. Each receiver front end provides low-noise impedance matching to the antenna/PCB trace impedance. The receiver channels are capable of supporting higher modulation schemes including M-QAM (Quadrature Amplitude Modulation) and OFDM (Orthogonal Frequency-Division Multiple Access). To process the higher modulated signals, the receivers down-convert the incoming RF signals to the baseband while extracting the in-phase (I) and quadrature-phase (Q) components. Beamforming of the four channels is performed in the baseband after I/Q extraction using a discrete-time beamformer [46, 47]. Time-delayed sampling of the extracted I/Q baseband signals, along with the LO (Local Oscillator) phase shifter, captures coherent samples from each element and enables wideband beamforming as shown in Fig. 4a [48]. Independent gain, phase, and delay tuning is provided for each of the I and Q channels to reduce the respective mismatches in each channel. A high-frequency external clock (2×2$\times$2 ×RF input frequency) is provided for the signal down-conversion and time-delayed sampling, which is susceptible to impedance variations due to deformation and printing imperfections. A tunable impedance matching network is used at the clock input to alleviate these impedance imperfections.
@@
A full break-down of the inks used in each layer are also shown in Fig. 2. The implemented BFIC is on layer 5 with Cu ink printed non-RF traces on layer 4. Cu clad is used for RF traces on layers 3 and antenna reflector on layer 2. Finally, Ag (silver) ink is used to print antennas on layer 1 (shown in Fig. 6b). Ag vias and Cu Ink vias that was fixed with Ag are used between the layers.
Figure 4: a. 4 element self-calibrated beamforming receiver architecture with LNTA, downconversion mixer, phase shifter, TIA and discrete time beamformer with calibration loop in feedback. The beamformer output is digitized using an off-chip ADC and fed back to the on-chip calibration loop. b.Self-calibration loop structure correction deformities of a large array c. Illustration of self-calibration loop finding optimum phase control code under dynamic deformation.
@@ 2.3 On-chip dynamic deformation correction
The BFIC output is fed into an integrated self-calibrating loop considering the fact that a beamformer has only one maximum point which corresponds to the main lobe. The side lobes are considered as local maxima. For each angle, corresponding to each deformation condition (single- or multiple-curvature), there is a unique phase shift combination (and corresponding phase code word) for each element. For an 8-bit phase shifter, there are thus $2^{8}$ possible phase shifts. At a given AoA and a given deformation (single- or multiple-curvature), for a $2\times 2$ array, there are $2^{8\times 3}$ phase-shifting options, where one element is taken as reference (labeled as REF receiver in Fig. 4a). Only one combination amongst this gives the maximum Signal-to-Noise Ratio (SNR) when the beam is perfectly aligned with the transmitter. Incremental time-based search counters to validate each option is impractical for aforementioned applications-on-the-move. A data-driven or model-driven ML model can provide the optimum combination quickly but gets easily stuck into an unknown condition considering the plethora of uncertainties in real applications. The proposed integrated loop ensures fast convergence to the optimum phase shifter combination.
@@ 3.2 IC design details
As mentioned in Section 2.2, the beamformer is realized using a discrete-time sample-and-hold architecture. The beamforming network consists of time-interleaved switched-capacitor array that samples signals from each channel. Based on the AoA and the delay experienced by each element, the sampling instance can be adjusted to extract coherent samples from each channel. This time delayed sampling is enabled by the phase-shifter (PS) in the LO path with the phase-interpolator (PI) in the sample-and-hold clock path which provides true-time-delay and enables beamforming over a wide fractional bandwidth without beam-squint issue [48]. Th 8-bit PS, designed using an inverter-based vector modulator architecture, offers $5.6^{\circ}$ phase resolution in each of the four quadrants. The inverter-based PI in the sampler clock provides 6-bit tunability with a delay resolution of 76 ps. This work considers deformation correction over $\approx$140 MHz bandwidth which requires the PS tuning only for correcting the beam-pointing error. Both PI and PS are designed in the clock path away from the signal path, to minimize direct noise injection into the signal.
@@
The captured coherent samples are integrated over a clock period for beamforming. An two-stage internally compensated OTA with 685MHz unity-gain bandwidth is designed for the beamforming. A 4.2 GHz external clock feeds the on-chip LO and sampler clock generator with a 4-bit tunable matching network as mentioned in Section 2.2. The clock generator includes static clock dividers, clock drivers, and regenerator circuits to minimize power consumption. The prototype BFIC occupies 2.56 $\text{mm}^{2}$ (Fig. 6a) area (including pads) and consumes 18.5 mW per channel.
@@
Independent self-calibration loops are provided for each of the phase shifters $\phi_{1}$, $\phi_{2}$, $\phi_{3}$, implemented in an all-digital fashion that facilitates fast synthesis. The perturbation frequency, $\omega_{\text{p}}$, is set to a low frequency of 30 rad/s generated using a LUT with 128 entries. Each entry in the LUT is 17-bit long with the first two bits representing sign and magnitude. The HPF cut-off is set as 5 rad/s, to capture the slow variations in the $BF_{\text{out}}$. The accumulator gain of each loop is set as follows: $a_{\phi\text{1}}=15$ , $a_{\phi\text{2}}=20$, and $a_{\phi\text{3}}=25$ to achieve faster convergence. Each loop provides a 16-bit digital output representing the phase code word for the corresponding phase shift control. The $BF_{\text{out}}$ captured from BFIC is normalized to a 16-bit input ranging from -31 to +32, where 5-bits represent the integer value, 1-bit represents the sign, and 10-bit for the fractional points. The calibration loop occupies 0.026 $\text{mm}^{2}$ area per PS while consuming 1.5 mW.
@@
Figure 5: a. Measurement setup for deformation correction verification and radiation pattern.b. Calibration loop reduces deformation error to $<1.5^{\circ}$ from $7^{\circ}$; corresponding change in phase shifter codes $\phi_{1}$, $\phi_{2}$, $\phi_{3}$ are captured c. Worst case deformation w/ 38cm radius by tiling four 2x2 conformal arrays d. Overall deformation correction test setup. e. Radiation pattern after deformation correction at 0° Azimuth.

### v3

@@ Abstract
Conformal phased arrays provide multiple degrees of freedom to the scan angle, which is typically limited by antenna aperture in rigid arrays. Silicon-based RF signal processing offers reliable, re-configurable, multi-functional, and compact control for conformal phased arrays that can be used for on-the-move communication. While the lightweight, compactness, and shape-changing properties of the conformal phased arrays are attractive, these features result in dynamic deformation of the array during motion, leading to significant dynamic beam pointing errors. We propose a silicon-based, compact, reconfigurable solution to self-correct these dynamic deformation-induced beam pointing errors. Furthermore, additive printing is leveraged to enhance the flexibility of the conformal phased arrays, as the printed conductive ink is more flexible than bulk copper and can be easily deposited on flexible sheets using different printing tools, providing an environmentally-friendly solution for large-scale production. The inks such as conventional silver inks are expensive and copper-based printable inks suffer from spontaneous metal oxidation that alters trace impedance and degrades beamforming performance. This work uses a low-cost molecular copper decomposition ink with reliable RF properties at different temperature and strain to print the proposed intelligent conformal phased array operating at 2.1 GHz. Proof-of-concept prototype $2\times 2$ array self-corrects the deformation induced beampointing error with an error $<1.25^{\circ}$. The silicon-based array processing part occupies only 2.58 mm2 area and 83 mW power per tile.
@@
Additive Printing, conformal phased arrays, deformation correction, beam stabilization, molecular copper decomposition ink, self-correction loop, scalable tile-based.
@@ 1 Introduction
Phased array systems support directional data transfer, which enables inherent spatial selection to support ever-increasing wireless data rate. Initial research on the phased arrays started several decades back (earliest reported in 1909) to enable inertia-free fast beam steering by avoiding mechanical gimballing [1, 2]. The underlying assumption in these works was that the antenna surfaces were rigid. By the 1940s, conformal arrays were developed to curve around spherical and nonuniform surfaces[3, 4, 5, 6] that afforded array systems an added degree of freedom by eliminating the conventional radome. Conformal phased arrays have been used extensively since then in the automotive, aviation, and space industries. Having a non-uniform antenna structure enhances the field of view (FoV). When deployed on an aircraft, the array benefits from the aerodynamic shape of the plane, which improves the fidelity of the in-flight wireless and makes the airborne systems such as Unmanned Aerial Vehicles (UAV) management easy, as shown in Fig. 1.
@@
Satellite communication (SATCOM) has embraced conformal phased arrays due to their lightweight, small volume, and shape-changing properties. The manageable costs associated with launching and deploying [7, 8] have led to the widespread adoption of conformal phased arrays in Low-Earth Orbit (LEO) satellites[9]. LEO satellite networks such as Amazon Kuiper[10], SpaceX Starlink [11], OneWeb, and Telesat Light-speed, boast coverage fields that collectively span the entire terrestrial surface area, opening profound opportunities for direct-to-consumer wireless communications in remote areas [12, 13]. These satellite networks, in conjunction with wearable/textile arrays[14, 15, 16], will further aid challenging research expeditions and facilitate disaster relief management as illustrated in Fig. 1.
Figure 1: Dynamic wing load variations and wing vibrations lead to deformed antenna arrays impacting in-flight wireless and navigation in Unmanned Aerial Vehicles (UAV); dynamic deformation in SATCOM and lightweight textile arrays. Silicon-based self-correction processor addresses gain loss and beam pointing error due to dynamic deformation in conformal phased arrays in real-time with the proposed additively printed, lightweight, tile-based flexible array.
@@
Surface deformation of a conformal phased array with curvature enhances its Field-of-View (FoV) but incurs increased inter-element coupling [17] and path length variations leading to beam pointing errors [18]. For satellites, it is possible to estimate the amount of deformation following the expansion of arrays once they are placed in orbit. While the initial deformation due to the aerodynamic shape of aircraft and drones can be estimated, the dynamic deformations arising from wing loading and vibrations[19] during transit cannot be predetermined. These dynamic deformations depend on factors such as array weight, size, drone/aircraft speed, wind conditions, initial deformation, etc. Intelligent self-adaptive conformal phased arrays will serve as the cornerstone to enable existing and emerging applications, as shown in Fig. 1, while overcoming limitations from dynamic deformations [20].
@@
Recent studies have explored numerous strategies for correcting the radiation pattern of conformal phased arrays [21, 22, 23] arising from physical deformities. However, element-level deformation sensing and compensation with conventional mechanical strain sensors [24, 25] and sensing mutual coupling [26, 27] has been ineffective due to complexity and non-planar shapes beyond single-point curvature. Machine-learning (ML), and in particular, pre-trained deep-learning-based networks, are increasingly being used for adaptive beam synthesis, allowing arbitrary shape correction based on training[19]. However, it relies heavily on feature-rich longitudinal-series training data with prior knowledge of the deformation surface and uncertain operational environments. These models further require separate dictionary sets for near-field and far-field patterns as well as environmental conditions, demands exponential increase in storage as the array size is scaled. The multi-dimensional variable dynamic range search algorithm in [28], employs array-level compensation suited for both near- and far-field. However, its higher computational complexity and the additional requirement for separate generation and recovery units with high computing power along with low-power transceivers make this technique challenging to adapt for airborne platoons and portable units. Phase correction using iterative methods such as genetic algorithm [29] and iterative phase synthesis [30, 31, 32] not only requires external engines for executing the iterative algorithm, but also incurs large latencies not suited for fast communications-on-the-move applications.
@@
Figure 2: Tile based array scaling with edge computing for gain, phase, delay and impedance self-calibration loops (conceptual). The proposed additively printed $2\times 2$ conformal tile layer stackup with beamforming IC and a. Full break down of Inks in each layers of array, Ball Milled Cu Ink b. SEM (Scanning Electron Microscope) of the ink before sintering, and after sintering c. Test setup for resistance measurements at different temperatures: Printed traces using the copper molecular decomposition ink are placed on a hot plate, and the temperature is varied to note the variation in resistance with respect to temperature. d. Measured resistance at different temperatures.
@@ 2.1 Low-cost reliable molecular copper decomposition ink for additive printing
Additively printed RF circuits have gained prominence in next-generation wireless systems with their ability to generate non-uniform conformal shapes required for large-scale infrastructure while being environmentally-friendly [33, 34, 35, 14, 15, 16]. Recent works for additive printing techniques such as aerosol jetting, ink jetting, direct writing, and screen printing have shown enhanced scalability, reduced material waste generation, and cost-effective manufacturing processes. However, the efficacy of the additive printing method is heavily contingent upon the quality and properties of the conductive inks utilized in the application. Common challenges in additive manufacturing of conductive inks revolve around printability and electrical performance [36]. Factors such as additives, ink viscosity, uniformity, micro/nanomaterial structure, and size directly influence the potential printing method, substrate material, sintering conditions, and electrical performance of the traces. The electrical performance of printed conductive traces is also commonly evaluated for stability under repeated exposure to high temperatures, corrosive environments, and mechanical bending and torsion, reflecting long-term reliability during use.
@@
Figure 3: a. Dipole antenna printed using proposed Cu decomposition ink testing under varied strains b. Laser ablated dipole antenna tested at different temperatures on a hotplate c. Test set up for measuring ink stability under different strains. d. Measured S11 of the dipole antenna printed using Cu ink to demonstrate ink stability at different temperatures and e. strain.
Although silver (Ag)-based inks have been largely appreciated due to their low resistivity (1.59 $\mu\Omega.cm$) for use in printed electronics, their high-cost margins make copper (Cu) a much more cost-effective alternative with potential for comparable resistivity performance (1.72 $\mu\Omega.cm$) [38, 39]. However, excessive additives needed to improve printability of Cu nanostructured inks and with a greater potential for ambient condition oxidation suggests that molecular Cu-based inks are more advantageious for long term commercial and industrial applications [39, 40, 41]. Herein a molecular Cu formate-based ink is utilized to create a highly conductive Cu thin-film ($=$35 MS/m)[40] on a Pyralux substrate with demonstratively consistent electrical, and RF performance under varied strains and temperatures.
Figure 4: a. 4-element beamforming receiver architecture with four downconverters and discrete-time beamformer with self-corrected calibration loop in feedback. The beamformer output is digitized using an off-chip ADC and fed back to the on-chip calibration loop. b.A single tile self-calibration loop structure for correcting beam pointing error c. Single channel receiver architecture with LNTA, downconversion mixer, phase shifter, TIA, and time-delayed sampler. d.Illustration of self-calibration loop finding optimum phase control code under dynamic deformation.
@@ 2.2 Scalable tile-based delay-phase tunable receiver array
The emergence of highly integrated silicon-based RF signal processors has revolutionized array synthesis and beam steering with multi-modal spatial signal processing [42, 43, 44], enabling high Signal-to-Noise Ratio (SNR) in commercial arrays. When the number of elements increases, the beam width reduces, thereby enhancing the spatial selectivity. At 2.1 GHz, the antenna dimensions are in centimeters, and therefore, scaling to a larger array is not straightforward. This scaling comes with several challenges [45], such as additive printing of high-resolution structures with physically large dimensions and combining signals from multiple antennas over wide bandwidth with high signal-to-noise ratio. We are adopting a modular tile-based approach for the receiver array as shown conceptually in the illustration in Fig. 2. In this work, each tile consists of a $2\times 2$ conformal antenna array attached to a 4-channel beamforming integrated circuit (BFIC). Figure 5e shows four tiles thermally bonded together to form a 16-element array. The BFIC in each tile combines each antenna signal and, therefore, acts as a sub-array. Each sub-array outputs are then combined using a second-stage combiner to achieve $4\times 4$ radiation pattern. The gain, phase shift, and impedance mismatch, and delay skews exacerbates as the array scales. The modular tile-based array structure thus enables easier calibration of larger arrays. An N-element array can be divided and processed into N/4 sub-arrays.
@@
The BFIC in a $2\times 2$ tile performs discrete-time true-time-delay BFIC as shown in Fig. 4a. The developed ink and the printing technique ensure that the four channels of the BFIC are connected to each antenna through impedance-matched via holes and traces. Each receiver front end provides a low-noise impedance matching to the antenna/PCB trace impedance. The receiver channels are capable of supporting higher modulation schemes, including M-QAM (Quadrature Amplitude Modulation) and OFDM (Orthogonal Frequency-Division Multiple Access). To process the higher modulated signals, the receivers down-convert the incoming RF signals to the baseband while extracting the in-phase (I) and quadrature-phase (Q) components. Beamforming of the four channels is performed in the baseband after I/Q extraction using a discrete-time beamformer [46]. Time-delayed sampling of the extracted I/Q baseband signals, along with the LO (Local Oscillator) phase shifter, captures coherent samples from each element and enables wideband beamforming as shown in Fig. 4a [47]. Independent gain, phase, and delay tuning is provided for each of the I and Q channels to reduce the respective mismatches in each channel. A high-frequency external clock (2×2$\times$2 ×RF input frequency) is provided for the signal down-conversion and time-delayed sampling, which is susceptible to impedance variations due to deformation and printing imperfections. A tunable impedance matching network is used at the clock input to alleviate these impedance imperfections.
@@
A full breakdown of inks used in each layer are also shown in Fig. 2. The implemented BFIC is on layer 5 with Cu ink printed non-RF traces on layer 4. Cu clad is used for RF traces on layer 3 and antenna reflector on layer 2. Finally, Ag (silver) ink is used to print antennas on layer 1 (shown in Fig. 6e). Ag vias and Cu Ink vias that were fixed with Ag are used between the layers.
@@ 2.3 On-chip dynamic deformation correction
The BFIC output is fed into an integrated self-calibrating loop considering the fact that a beamformer has only one maximum point, which corresponds to the main lobe. The side lobes are considered as local maxima. For each angle, corresponding to each deformation condition (single- or multiple-curvature), there is a unique phase shift combination (and corresponding phase code word) for each element. For an 8-bit phase shifter, there are thus $2^{8}$ possible phase shifts. At a given AoA and a given deformation (single- or multiple-curvature), for a $2\times 2$ array, there are $2^{8\times 3}$ phase-shifting options, where one element is taken as reference (labeled as REF receiver in Fig. 4a). Only one combination among these gives the maximum SNR when the beam is perfectly aligned with the transmitter. Incremental time-based search counters to validate each option is impractical for aforementioned applications-on-the-move. A data-driven or model-driven ML model can provide the optimum combination quickly but gets easily stuck into an unknown condition considering the plethora of uncertainties in real applications. The proposed integrated loop ensures fast convergence to the optimum phase shifter combination.
@@ 3.2 IC design details
As mentioned in Section 2.2, the beamformer is realized using a discrete-time sample-and-hold architecture. The beamforming network consists of time-interleaved switched-capacitor array that samples signals from each channel. Based on the AoA and the delay experienced by each element, the sampling instance can be adjusted to extract coherent samples from each channel. This time delayed sampling is enabled by the phase-shifter (PS) in the LO path with the phase-interpolator (PI) in the sample-and-hold clock path which provides true-time-delay and enables beamforming over a wide fractional bandwidth without beam-squint issue[47]. Th 8-bit PS, designed using an inverter-based vector modulator architecture, offers $5.6^{\circ}$ phase resolution in each of the four quadrants. The inverter-based PI in the sampler clock provides 6-bit tunability with a delay resolution of 76 ps. This work considers deformation correction over $\approx$70 MHz base-band bandwidth which requires the PS tuning only for correcting the beam-pointing error. Both PI and PS are designed in the clock path away from the signal path, to minimize direct noise injection into the signal.
@@
The captured coherent samples are integrated over a clock period for beamforming. A two-stage internally compensated OTA with 685MHz unity-gain bandwidth is designed for the beamforming. A 4.2 GHz external clock feeds the on-chip LO and sampler clock generator with a 4-bit tunable matching network as mentioned in Section 2.2. The clock generator includes static clock dividers, clock drivers, and regenerator circuits to minimize power consumption. The prototype BFIC occupies 2.56 $\text{mm}^{2}$ (Fig. 6a) area (including pads) and consumes 18.5 mW per channel.
@@
Independent self-calibration loops are provided for each of the phase shifters $\phi_{1}$, $\phi_{2}$, $\phi_{3}$, implemented in an all-digital fashion that facilitates fast synthesis. The perturbation frequency, $\omega_{\text{p}}$, is set to a low frequency of 30 rad/s generated using a LUT with 128 entries. Each entry in the LUT is 17-bit long with the first two bits representing sign and magnitude. The HPF cut-off is set as 5 rad/s, to capture the slow variations in the $BF_{\text{out}}$. The accumulator gain of each loop is set as follows: $a_{\phi\text{1}}=15$ , $a_{\phi\text{2}}=20$, and $a_{\phi\text{3}}=25$ to achieve faster convergence. Each loop provides a 16-bit digital output representing the phase code word for the corresponding phase shift control. The $BF_{\text{out}}$ captured from BFIC is normalized to a 16-bit input ranging from -31 to +32, where MSB bit represents the sign, next 5-bits represent the integer value, and lower 10-bit for the fractional part. The calibration loop occupies 0.026 $\text{mm}^{2}$ area per PS while consuming 1.5 mW.
@@
Figure 5: Measurement setup for deformation correction verification and radiation pattern. a. Illustration of test setup. b. Test bench for over-the-air measurements. c. Calibration loop reduces deformation error to $<1.5^{\circ}$ from $7^{\circ}$; corresponding change in phase shifter codes $\phi_{1}$, $\phi_{2}$, $\phi_{3}$ are captured d. Radiation pattern without deformation and with deformation after correction at $0^{\circ}$ Azimuth. e. (left) weight measurement of $2\times 2$ tile with BFIC and its connections, (right) Measurement setup for worst-case deformation with 38 cm radius of curvature by tiling four $2\times 2$ conformal arrays.